OpenStudy (anonymous):

When finding the sum of the arithmetic series 4 + 1 + -2 + -5 + -8 + -11, you will have a step of work that contains the number -3. Use complete sentences to describe how the -3 was calculated and what it means.

jimthompson5910 (jim_thompson5910):

4+1 = 5 4+1-2 = 3 4+1-2-5=-2 4+1-2-5-8=-10 4+1-2-5-8-11=-21 So clearly there isn't a -3 in this whole process. Something seems off, but idk what.

OpenStudy (anonymous):

wait so do you just add all of the numbers to find the sum?

jimthompson5910 (jim_thompson5910):

yes because "sum" means "result after addition"

jimthompson5910 (jim_thompson5910):

I'm not sure where they're getting -3 though...

OpenStudy (anonymous):

ha i'm so confused?

jimthompson5910 (jim_thompson5910):

yeah same here, so there has to be a typo

jimthompson5910 (jim_thompson5910):

that's my guess

OpenStudy (anonymous):

okay, thank you!!

jimthompson5910 (jim_thompson5910):

np, sorry I couldn't be of more help

OpenStudy (anonymous):

that's okay, i appreciate it :)

jimthompson5910 (jim_thompson5910):

SmallChain, that's something for a private message and it's not really relevant.

OpenStudy (anonymous):

^truth, do you understand this smallchain??

OpenStudy (e.mccormick):

Know how to find the equation for an arithmetic series?

OpenStudy (anonymous):

i still don't really understand. do you think you could walk me through it please?

mathslover (mathslover):

Well I think , it is the common difference which is equal to (-3)

OpenStudy (e.mccormick):

Do you see the sequence of numbers? 4, 1, -2, -5, -8, -11 Yes. That is the part I was going to point out.

mathslover (mathslover):

And the step that involves the use of "-3" is : \(S_n = \cfrac{n}{2} ( 2a + (n-1)d ) \) where S_n = sum of n-terms n = no. of terms a = first term d = common difference.

mathslover (mathslover):

It is actually , Arithmetic Progression (as said earlier by @e.mccormick ) . @girlscum - Are you getting my point?

OpenStudy (e.mccormick):

You look at each step, and what number added to a step in the sequence would get you to the next step? That is the common difference. So you can find it using any two steps and check it with any two others. 4+d=1, solve for d. then -5 + (-3) = -8 to check.

OpenStudy (e.mccormick):

So write about that and you will have your answer.

OpenStudy (anonymous):

i think i understand.. how exactly are you finding the common difference though? sorry i'm terrible at these problems.

mathslover (mathslover):

You can find the common difference by subtracting the two consecutive terms : a_1 , a_2 , ... , a_n common diff. = a_2 - a_1 = a_3 - a_2 = .... = a_n - a_(n-1)

OpenStudy (anonymous):

so the answer would be -3 is the common difference?

mathslover (mathslover):

Not actually,you will have to explain how u calculated -3 and what is common difference.

OpenStudy (anonymous):

okay so you'd subtract consecutive terms to find the common difference which is -3?

mathslover (mathslover):

Yeah, it is -3 here.

mathslover (mathslover):

For example : Let us take a series : 2,4,6,8 common difference = 4-2 = 6-4 = 8 - 6 = 2

OpenStudy (anonymous):

okay i get it now thank you both so much! do you think you could help me with some other problems?

mathslover (mathslover):

Well, I think, I can! I will try my best :) Also, Since you are new to OpenStudy : Welcome to OpenStudy.

OpenStudy (anonymous):

thank you! hold on one sec let me find just get my work out

OpenStudy (anonymous):

What is the common difference of a 43-term arithmetic sequence where the first term is -13 and the sum is 9,374?

mathslover (mathslover):

Use the formula : \(S_n = \cfrac{n}{2} ( 2a + (n-1)d ) \)

OpenStudy (anonymous):

how would i find d?

OpenStudy (anonymous):

@mathslover

mathslover (mathslover):

S_n = 9,374 (given) n = 43 a = -13 put in the values and solve for d.

OpenStudy (anonymous):

oh, my bad i misread the problem.

OpenStudy (anonymous):

okay i think i did it completely wrong do you think you could walk me through it haha

OpenStudy (anonymous):

if it helps the answer has to be one of these 11 11.5 12 12.5

OpenStudy (anonymous):

@mathslover